Suppose that Kiara Corp’s stock just paid a dividend of $2, for which rs = 12% and g = 20% for 4 years before achieving long-run growth of 4%. What is the price of the stock today?

$35.74

None of these

$34.26

$41.42

$39.42

Beta

A measure of the idiosyncratic risk present in an asset.

A measure of a stock’s price.

A measure of the sensitivity of expected returns of a risky asset to movements in the market portfolio.

A measure of the extent to which two stocks are related.

None of these.
Answer #1

1)

Dividend today, D0 = $ 2

g = 20%

rs = 12%

D1 = 2* 1.2 = 2.4

D2= 2.4 * 1.2 = 2.88

D3 = 2.88 * 1.2 = 3.46

D4 = 3.456 * 1.2 = 4.15

Terminal value = 4.15 * (1+4%)/(12% - 4%) = 53.91

Price of the stock today = 2.4/(1.12) + 2.88/(1.12)^2 + 3.46/(1.12)^3 + 4.15/(1.12)^4 + 53.91/(1.12)^4

= 43. 79

Answer is None of these

2)

Beta is A measure of the sensitivity of expected returns of a risky asset to movements in the market portfolio. Beta measures the systematic or market risk of firm. This risk as it is macro cannot be diversified through traditional diversification strategy.